find F(4) for polynomial F(x)=2x^2-x+9
Answer by jim_thompson5910(35256)

F(x)=2x^2-x+9

F(4)=2(4)^2-4+9

F(4)=2(16)-4+9

F(4)=32-4+9

F(4)=37


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim